OpenStudy (quickstudent):

What is the 7th number in the pattern: 1, 4, 9, 16, 25, …? 32 36 39 49

OpenStudy (mathmale):

1, 4, 9, 16 and 25 can all be characterized by the word ...... ? Once you've spotted this pattern, finding the next number should be easy.

OpenStudy (whyy):

Still need help?

OpenStudy (whyy):

The pattern is going by odds, 1,4,9,16,25 just place in the missing numbers 1+3=4+5=9+7=16+9=25+11=36

OpenStudy (osprey):

looks like they're all PERFECT SQUARES to me. 1squared 2 squared 3 squared etc. So, a bit of times tables ...

OpenStudy (mathmale):

WHYY: That's an interesting observation. Hadn't noticed this myself. Osprey: You're right on target.